### v1

# An Survey of the Evolution of Language Model-Based Dialogue Systems

Abstract
Dialogue systems (DS), including the task-oriented dialogue system (TOD) and the open-domain dialogue system (ODD), have always been a fundamental task in natural language processing (NLP), allowing various applications in practice. Owing to sophisticated training and well-designed model architecture, language models (LM) are usually adopted as the necessary backbone to build the dialogue system. Consequently, every breakthrough in LM brings about a shift in learning paradigm and research attention within dialogue system, especially the appearance of pre-trained language models (PLMs) and large language models (LLMs). In this paper, we take a deep look at the history of the dialogue system, especially its special relationship with the advancements of language models. Specifically, we delineates the evolutionary trajectory of LM-based dialogue systems by categorizing it into four distinctive stages, each marked by pivotal breakthroughs in language models: 1) early stage, which witnesses predominantly statistical LMs, resulting in rule-based or machine-learning-driven DS; 2) the independent development of TOD and ODD based on neural language models (NLM; such as LSTM and GRU), since NLMs lack intrinsic knowledge in their parameters; 3) fusion between different types of dialogue systems with the advert of pre-trained language models (PLMs), starting from the fusion between four sub-tasks within TOD, and then TOD with ODD; and 4) current LLM-based dialogue system, wherein LLMs can be used to conduct TOD and ODD seamlessly. In this way, our survey offers a systematic perspective, categorizing the stages in a chronological order aligned with LM breakthroughs, providing a comprehensive review of state-of-the-art research outcomes. What’s more, we turn our attention to emerging topics and engage in a discussion on open challenges, providing valuable insights into the future directions for LLM-based dialogue systems. In summary, this survey delves into the dynamic interplay between language models and dialogue systems, unraveling the evolutionary path of this essential relationship. Through this exploration, we pave the way for a deeper comprehension of the field, guiding future developments in LM-based dialogue systems.
Index Terms:
Task-oriented Dialogue System, Open-domain Dialogue System, Pre-trained Language Models, Large Language Models.
1 Introduction
Building a conversational intelligent system has always been a fundamental objective in the realm of natural language processing [1, 2]. The ability to interact naturally and seamlessly with machines has opened up new avenues for human-machine communication, leading to transformative applications across different industries, from Siri, Xiaoice [3] to the latest New Bing  [^1] https://www.bing.com/new and Google Bard  [^2] https://bard.google.com/. At the heart of this evolution is the pivotal role that dialogue systems play, which are designed to facilitate interactions characterized by harmlessness [4], helpfulness [5], trustworthiness [6], and personalization [7]. These systems aim to emulate human-to-human conversations, thereby offering the potential to enhance user experiences, streamline tasks, and provide personalized assistance across domains such as customer support, virtual assistants, healthcare, education, and more, depending on the types of dialogue context [8] as shown in Table I.
Specifically, there are two major types of dialogue in practice: task-oriented dialogue and chit-chat. Task-oriented dialogue revolves around assisting users in achieving specific tasks or goals, such as making reservations or booking tickets. In this context, the dialogue system acts more like a helpful assistant, providing relevant information and guidance to users [9, 2]. Conversely, chit-chat involves casual and informal conversations, mainly aimed at establishing and maintaining social connections. Here, the dialogue system adopts the role of a friendly chatbot, engaging users in a more relaxed and conversational manner [1]. Based on their distinct roles, two types of dialogue systems have been proposed: Task-oriented Dialogue System (TOD) and Open-domain Dialogue System (ODD). The TOD is designed to efficiently handle task-oriented conversations, guiding users towards accomplishing specific objectives by detecting user intentions, tracking dialogue state, making suitable actions, and responding accordingly. In contrast, the ODD is intended for open-domain interactions, enabling free-flowing conversations on a wide range of topics by directly mapping the dialogue context into the response, without a predefined task or goal. These two types cater to different use cases and user needs, showcasing the versatility and applicability of dialogue systems in diverse scenarios. Thus, most of the previous works (before $\sim$2020, Figure 2) aim to design and build these dialogue systems independently using different backbones considering their different roles [2, 1].
Figure 1: The development of LM-based Dialogue Systems, which can be divided into four distinct stages, each marked by significant breakthroughs that have paved the way for advancements in the field: 1) Early Stage (1966 - 2015); 2) The Independent Development of TOD and ODD (2015 - 2019); 3) Fusions of Dialogue Systems (2019 - 2022); 4) LLM-based DS (2022 - Now).
However, the revolutionary progress in Language Models (LMs) has sparked a transformative evolution in dialogue systems, reshaping their very foundation. The core objective of LMs is to predict word sequence probabilities has brought about a profound impact on dialogue systems [10, 11], empowering them to learn world knowledge from the pre-training corpus [12] and generate more contextually relevant and helpful responses [13]. Concurrently, each milestone in LM development has acted as a catalyst for transformative changes in dialogue systems, playing a pivotal role in their rapid evolution and ongoing trends, especially the appearance of latest large language models (LLMs), such as ChatGPT. Rather than dwelling on the development of language models, our focus lies in the enthralling journey of dialogue systems undergoing this transformative process, which can be divided into four major development stages corresponding to the four stages of LM [14] as shown in Figure 1:
TABLE I: Different types of dialogues in TOD and ODD. We keep the last turn of users for conciseness and brevity.
[…]
  • Early stage – Statistical language models (SLMs). The first dialogue system — Eliza, was proposed by MIT in 1966, which was even earlier than the rise of SLMs in the 1990s, allowing plausible conversations between humans and machines. Subsequently, a series of dialogue systems were constructed, including a travel agent aiding the client to make a simple return trip to a single city in California [15] and a rule-based chatbot configured with a persona, employing heuristic pattern matching rules to address inquiries about ages, hobbies, interests, and etc [16]  [^3] https://web.njit.edu/~ronkowit/eliza.html. In addition to academic endeavors, more and more companies pay attention to the field of dialogue systems and have developed their own products. For example, IBM has developed WATSON, Apple has Siri, and Microsoft has XiaoIce. These early virtual assistants have limited functionalities and follow a rigid process, primarily assisting users with single and simple tasks like scheduling appointments, setting reminders, and providing basic answers to questions. At this stage, most of these dialogue systems are task-oriented, the prevailing nature of these dialogue systems is predominantly task-oriented, employing either modular-based approaches or machine-learning methods based on SLMs.
  • The independent development of TOD and ODDNeural language models (NLMs). Around 2015, there was a sea change when the sequence to sequence (seq2seq) framework was proposed, in which salient features were learned jointly with the training of the model — NLM [17, 18]. Specifically, NLMs characterize the probability of word sequences by neural networks, e.g., two typical recurrent neural networks (RNNs) such as long short-term memory networks (LSTM) [19] and gated recurrent units (GRU) [20]. These NLMs serve as the backbone to map variable-length input sequences to variable-length output sequences which is achieved by employing two main components: an encoder and a decoder. This paradigm shift facilitated the emergence of Open-domain Dialogue (ODD) systems ($\sim$2015, Figure  2), garnering attention due to a well-defined end-to-end framework enabled by the seq2seq network. Simultaneously, NLMs were also applied to TOD, with the goal of achieving better performance and performing complex tasks in multiple domains [21, 22, 23].
  • Fusions of Dialogue Systemspre-trained language models (PLMs). Traditional NLMs requires extensive hand-crafted features selection and domain-specific knowledge, making them cumbersome and time-consuming to develop for each task. However, with the advent of deep learning and large-scale language corpora, pre-trained language models (PLMs) emerged as a groundbreaking solution. These models are pre-trained on massive amounts of unlabeled text data, capturing rich semantic and syntactic patterns [24, 11, 25]. By leveraging pre-training, subsequent fine-tuning on specific tasks becomes more accessible, enabling the models to adapt quickly and achieve impressive results with minimal task-specific data. In the realm of dialogue systems, DialoGPT exemplifies this approach by undergoing pre-training on 147 million conversation-like exchanges extracted from Reddit comment chains and then finetuning on a few training examples to generate more relevant, informative, and context-consistent responses [13]. This paradigm, commonly known as pre-train and finetune [26], leads to the resulting model as pre-trained dialogue model (PDM) that undergoes fine-tuning using a dialogue corpus. It is worth noting that the major differences between PDMs lies in the the corpus used for training and the backbone language models adopted, such as PLATO [27, 28], BlenderBot [29], Meena [30], Pangu-bot [31], PLUG [32], and etc. In addition, there is an increasing focus on the fusion of different tasks inside TOD (e.g., End-2-end TOD), TOD with ODD, and propose Unified Dialogue System (UniDS) in this period (Figure 2).
  • LLM-based dialogue systemLarge language models (LLM). Recognizing the correlation between increased pre-training corpus size and model size with enhanced performance across various NLP tasks [33], researchers try to scale the model size and the size of pre-training corpus simultaneously to be more sample efficient, enabling the model to learn more intricate patterns and representations from the data. Thus, the PLMs become Large Language Models (LLMs), such as GLM [34], LLaMA [35], and InstructGPT [4]. Due to the larger model size and the utilization of extensice, high-quality pre-training corpora, these LLMs offer unprecedented capabilities in many language understanding and generation tasks, including but not limited to question-answering and named entity recognition, thus reshaping the landscape of dialogue systems. With further instruction-tuning using data as (instruction, input, output), these LLMs are capable of answering diverse questions and following various instructions such as writing an email and telling a joke. At this moment, the LLMs can be directly used as the dialogue systems. Moreover, finetuning the LLM with in-domain data, i.e., dialogue/conversational corpus, can further enhance its conversational ability. For example, the evolution of GPT-3.5 to InstructGPT, and then ChatGPT exactly follows this path. Similarly, ChatGLM [36] is also finetuned using publicly available Chinese dialogue datasets based on the corresponding LLM – GLM. To enhance alignment with human preferences and values, the incorporation of reinforcement learning becomes imperative, particularly in the final stages of model refinement. This involves leveraging feedback from either human evaluators [4] or automated systems [37] to iteratively improve the model’s performance and ensure its adherence to desired criteria. After all, this type of LLM-based dialogue system (LLM-based DS) performs well at both task-oriented dialogue and chit chat, making it an ideal foundation for a general conversational AI.
It is worth noting that the development of LM-based Dialogue System is a ongoing and evolving process, and the stages are not rigidly separated by specific turning points. Instead, there are overlaps and continuous advancements across these stages, particularly in the later stages. Generally, with the advancement of LM, the boundary between TOD and ODD and the boundary between DM and LM have become increasingly blurred, ushering in a new era of LLM-based Dialogue System. Despite that the existing studies have thoroughly examined the development and challenges associated with different types of dialogue systems, such as TOD [2] and ODD [1], there still remains a notable gap in the literature when it comes to offering a holistic perspective on the evolutionary trajectory of dialogue systems, particularly in light of the advancements in LM. By critically analyzing and understanding the impact of pivotal milestones, we can gain valuable insights into the possibilities that await and the transformative impact dialogue systems may have on diverse domains and applications in the foreseeable future. This survey aims to present a comprehensive overview of the realm of LM-based dialogue systems  [^4] We have made thorough efforts to cover a wide range of works, but acknowledge the possibility of omissions. Readers are encouraged to provide suggestions for any missed contributions or errors. We are committed to updating the article with new approaches or definitions proposed by the community., shedding light on the prominent directions which the field is currently progressing. This entails a meticulous review of the existing literature, aimed at elucidating the paradigm shifts observed within different stages of LM-based dialogue systems. These shifts encompass a range of integrations across various levels, notably encompassing the task, data, and framework dimensions.
Figure 2: The number of publication at ACL Anthology conferences of different research directions in LM-based dialogue systems (The raw data is downloaded around 31/10/2023).
The remainder of this survey is organized as follows: Section 2 provides the description of different types of dialogues and define the problem in a unified manner, followed by the early development of two types of LM-based dialogue systems: TOD and ODD in Section 3. Sections  4 review and summarize the evolution of different dialogue systems when PLMs are scaled to LLMs. We then investigate emerging trends and open challenges in LLM-based DS (See Section 5). Then, Section 6 discusses the open problems for LLM-based dialogue systems. Finally, we conclude the survey in Section 7 by summarizing the major findings and discussing the remaining issues for future work.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hiddendraw, align=left, minimum width=2.5em, s sep=6pt, inner xsep=2pt, inner ysep=1pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=2em,font=, where level=2text width=6.5em,font=, where level=3text width=16em,font=, where level=4text width=16em,font=, [LM-based
Dialogue System [TOD [Natural Language
Understanding (NLU) [RNN: RNN-LU [38, 39]
LSTM: BiLSTM [40, 41], Attention-BiLSTM [42]
Slot-Gated [43] , DF-Net [44] [BERT: Joint-NLU, [45] BERT-Joint [46]
StackNLU [47]
RoBERTa: MultiDomain [48]
Others: MinDS [49] ] ] ] [Dialogue State
Tracking (DST) [RNN: Word-level RNN [50], NBT [51], MD-DST
[52], MemN2N [53], GLAD [54]
GRU: TRADE [55], SAS [56] ,
LSTM: FJST [57], [Bert: CSFN [58], SUMBT [59], EMD [60],
DS-DST [61]
Albert: DiCoS [62]
GPT-2: NP-DST [63], SimpleToD [9],
SOLOIST [64] ] ] ] [Dialogue Policy
Learning (DPL) [RNN: LRU [65], RS-RNN [66], MADPL [67]
LSTM: OnlineRL [68], GDPL [69], AdvL [70] [BERT: PPO-OFF-Comb [37]
GPT2: SimpleToD [9] ] ] ] [Natural Language
Generation (NLG) [ LSTM: SC-LSTM [71], Meta-NLG [72], HDSA [73] [GPT2: SimpleToD [9], SC-GPT [74]] ] ] [End-to-End (E2E) [LSTM: Network [22] [See Section 4.1] ] ] ] [ODD [ Generative DS [ RNN: NRM [75], Ext-ED [76]
LSTM: NCM [77], P-NCM [78], PersonaChat [79]
GRU: CoPerHED [80], STAR-BTM [81],
CKG-aware [82], K-NCM [83]
] ] [ Retrieval-based DS [ RNN: SMN [95]
LSTM: DL2P [96], Match-LSTM [97],
Attentive-LSTM [98]
GRU: Multi-view [99], DUA [100],
CNN: ARC [101], RSTP [102], LSM [103]
] ] [ Hybrid [ RNN: XiaoIce [112], EnsembleGAN [113]
GRU: Ensemble [114], PTE [115]
LSTM: EED [116], RR [117], HybridNCM [118]
Figure 3: Taxonomy of LM-based Dialogue System, with representative LMs of each task.
2 Background
2.1 Types of Dialogues
Before we proceed, it is essential to figure out how many different types of dialogs are in our daily life. Table I demonstrates some examples of four major types of dialogs, including:
  1. (a)
    social chit-chat. This type of dialog involves casual and informal conversations aimed at establishing and maintaining social connections. It includes light-hearted discussions about topics such as current events [121, 122], personal experiences [79, 123], hobbies, and interests [124]. Social chit-chat is typically used to create a friendly atmosphere and build social bonds, requiring the DS to generate persona-consistent, empathetic, and compassionate responses besides helpful.
  2. (b)
    knowledge-grounded chit-chat. Knowledge-grounded dialogs revolve around exchanging information and discussing topics that require a certain level of expertise or specific knowledge. It can involve discussing academic subjects, professional domains, or specialized areas of interest. This type of dialog often focuses on sharing and acquiring knowledge [105, 125]. The knowledge can be formed in different formats such as graphs [126], tables [127], figures [128], or even a combination of them [129].
  3. (c)
    question answering. Question-answering dialogs involve a user asking specific questions, and the dialogue system providing relevant and accurate answers. These dialogs are typically focused on obtaining specific information or clarifications. The system’s role is to understand the user’s queries and provide appropriate responses based on the available knowledge or resources [122, 130].
  4. (d)
    task-oriented dialog. Task-oriented dialogs center around assisting users in accomplishing specific tasks or goals. Users seek help from the dialogue system to perform actions such as making reservations, booking flights, ordering products, or seeking recommendations. These dialogs often involve a series of steps or sub-tasks that need to be completed to achieve the overall objective [21, 131, 132].
These are general descriptions of the different types of dialogs, and there may be variations or combinations of these types in real-world dialog scenarios [8]. The diversity and complexity of each type of dialog not only contribute to the wide range of applications for dialogue systems but also give rise to various challenges and problems. These challenges pose unique requirements and considerations, necessitating the development of dialogue systems that are both versatile and robust that can effectively handle the intricacies of each dialog scenario in a unified manner.
2.2 Task Definition
Following the unified definition recently [8, 133], we formulate the task of dialogue system as a seq2seq problem by directly mapping the dialogue context to the final response with or without the assistance of different external sources such as the local database in TOD, different knowledge bases in ODD, e.g., persona and documents. Formally, given the dialogue context $c=\{u_{1},s_{1},...,u_{t}\}$, and optionally the external knowledge bases $K$, the goal of the dialogue system is to generate a helpful and harmless response $r$ as shown below:
$r=\boldsymbol{M}(c,K)\quad or\quad\boldsymbol{M}(c)$ (1)
Where $\boldsymbol{M}$ is parameterized by different language models. Specifically, there are significant differences in interactions between various dialogue systems with the external knowledge base $K$. In this work, we discuss two major interactions corresponding to TOD and ODD respectively.
Formally, the TOD first generates the belief state according to the given context [134, 64]:
$b=\boldsymbol{M}(c)$ (2)
and then use it as a query to search related information in the local database $K$. The returned result $s$ is used as additional information to generate the final response, the entire process can be defined as follows:
$TOD(r|c,K)=f(r|c,b,s)f(b|c)$ (3)
Secondly, for the ODD, the processing is relatively simpler by directly rewriting the dialogue context as the query or using the whole dialogue context to search the external knowledge base $K$ [105]. In this case, TOD can be regarded as one special type of ODD with the source of knowledge as corresponding database [135].
$ODD(r|c,K)=f(r|c,q,s)f(q|c)$ (4)
It is important to clarify here that ODD contains the cases which does not require any external knowledge, by directly mapping dialogue context to corresponding responses using LMs [136].
3 The Independent Development of TOD & ODD
3.1 The Development of TOD
In the nascent stages of task-oriented dialogue system development, exemplified by systems like GUS [15] and ALICE [16], the technology relied heavily on rule-based approaches. These early systems operated within single domains and adhered to pre-defined decision trees or rule sets. For example, GUS was constrained to specific domains such as flight information retrieval [15], while ALICE was designed to simulate natural language conversation, following a script-like structure based on pattern matching and predefined responses [16]. These systems showcase impressive capabilities within their limited contexts but struggle to handle nuanced or complex conversations. They often require extensive manual crafting of rules and lack the ability to generalize beyond their narrow domains, making them ill-equipped for real-world applications with varying user intents and open-ended dialogues.
With the advent of the neural language model (NLM) and subsequent advancements in recurrent neural networks (RNNs), the landscape of task-oriented dialogue systems underwent a first evolution. The introduction of RNNs allows for more flexible and adaptive responses, enabling systems to process and generate natural language in a contextually aware manner. Dialogue systems now exhibit a broader spectrum of capabilities, embracing multi-domain functionality and accommodating diverse user intents. Unlike their predecessors, these modern systems can handle complex, intertwined tasks within a single conversation. For instance, a task-oriented dialogue system might seamlessly assist a user in booking a flight, reserving a hotel room, and planning an itinerary, all within the same interaction. This expanded capability has been made possible by the integration of four different components: Natural Language Understanding (NLU) [137], Dialogue State Tracking (DST) [138], Dialogue Policy Learning (DPL) [139], and Natural Language Generation (NLG) [140, 141]  [^5] https://github.com/yizhen20133868/Awesome-TOD-NLG-Survey.git.
  • [leftmargin=*,topsep=0pt,itemsep=0pt,parsep=0pt]
  • Natural Language Understanding (NLU): NLU processes user utterances and extracts relevant semantic frames, consisting of two sub-tasks: intent classification and slot filling. NLU can be rule-based, where predefined patterns or rules are used to extract information, or it can employ deep learning techniques like intent classification and named entity recognition.
  • Dialogue State Tracking (DST): DST maintains and updates the state of the conversation as it progresses. It keeps track of the user’s goals, preferences, and any relevant context obtained from the dialogue context. DST can use a rule-based approach to update according to the results from the NLU module, or replace the NLU module and directly track the dialogue state from a dialogue context with similar architecture as NLU [142, 60].
  • Dialogue Policy Learning (DPL): The dialogue policy component determines the system’s actions or responses based on the current dialogue state. It maps the dialogue state to appropriate system actions, which can include asking clarification questions, providing recommendations, or executing relevant API calls. Dialogue policies can be rule-based, handcrafted, or learned using techniques like reinforcement learning [21, 69, 67, 143, 37] or supervised learning [144, 75]. More details can be found in the survey paper [139].
  • Natural Language Generation (NLG): NLG generates system responses in natural language based on the dialogue policy’s output. It transforms structured information or system actions into coherent and fluent sentences that can be understood by users. NLG can employ template-based approaches, rule-based generation, or more advanced techniques like neural network-based generation models [145, 146].
It is worth noting that most studies at this stage independently target different components and propose task-specific designs. Additionally, there are some works try to combine consecutive components, such as word-level DST (coupling NLU and DST) [50, 51, 52, 53] and word-level policy (coupling Policy and NLG). As shown in Figure 3, the evolutionary trajectory of each sub-module in TOD has been underpinned by the rapid progress of neural language models. Notably, a pivotal transformation has been observed in the transition from conventional neural language models (NLMs) to the more versatile pre-trained language models (PLMs). This shift has significantly contributed to the enhancement of the dialogue system’s capabilities and performance.
Moreover, there are two obvious trends in this progress. First, TOD initially is designed to perform specific tasks within a predefined domain, however, as the PLMs emerged and gained traction, this paradigm shifted towards pre-train and then finetune [26]. By taking advantage of intricate patterns and world knowledge in the parameters of PLMS, the following method based on PLMs can generalize across tasks [21] and domains [41], even languages [147]. As tasks increase in complexity, they give rise to emerging challenges, including but not limited to insufficient availability of annotated data and the exacerbation of error propagation phenomena. To address these issues, inspired by the notable achievements of PLMs, a considerable body of research endeavors to employ an akin approach to conduct pre-training or in-domain finetuning on corpora specific to these sub-tasks in TOD simultaneously, constituting the second discernible trend. For example, SimpleTOD [9] uses a single, causal language model trained on all sub-tasks recast as a single sequence prediction problem, leading to better performance at all sub-tasks and robustness to noises. Consequently, these obstacles between distinct components and tasks gradually diminish, and more unified end-to-end TODs are on the way (§ 4.1).
3.2 The Development of ODD
Distinguishing from TOD, open-domain dialogue systems (ODDs) aim to perform chit-chat with users without task and domain restrictions and are usually fully data-driven in an end2end manner [136, 1]. In 2016, soon after the emergence of the seq2seq framework, the Alexa Price challenge was proposed to advance the development of ODD that is able to converse coherently and engagingly with users on open-end topics such as sports, politics, and entertainment [148]. Generally, there are three types of open-domain dialogue systems: generative dialogue systems [75, 78, 81, 93, 149], retrieval-based dialogue systems [95, 98, 104, 107, 111], and hybrid dialogue systems [112, 113, 120]. In detail, the retrieval-based systems retrieve predetermined responses from a fixed set in response to user input. In contrast, generative systems construct responses from scratch using language models, resulting in diverse and contextually appropriate outputs. Hybrid systems combine both approaches, using retrieval for efficiency and generative methods for flexibility, aiming to achieve a balance between accuracy and creativity. The choice of system type depends on factors such as the application’s requirements, available resources, and the desired user experience. Some representative works are summarized in Figure 3.
Similar to the TOD, ODD also went through a shift from NLM to PLM since the appearance of GPT [11], which operates in a decoder-only framework and is trained by predicting the next token auto-regressively. Most ODD based on NLM utilizes two typical recurrent neural networks: LSTM [19] and GRU [20], such as response selection tasks [111], knowledge-grounded dialogue [76, 83, 150], persona-consistent dialogue [78, 80, 79, 151], keywords [82] and topic [81] guided dialogue. A similar trend continues except the backbone of the dialogue system becomes PLMs after the appearance of BERT [24] and GPT2 [25]. Inspired by the exceptional capability learned during the pre-training stage, lots of works either directly pre-train the model from scratch using the dialogue dataset or further conduct in-domain finetuning using the current PLM as the initial checkpoint. The model after pretraining on tremendous dialogue datasets is formally defined as pre-trained dialogue model (PDM).
Distinguish from TOD which usually adopts encoder-only PLM, there are usually two specific architectures of different PDMs: encoder-decoder [30, 85, 29] and decoder-only [13, 94], especially for generative and hybrid ODD. The distinction here can be attributed to the core nature and objectives of TOD and ODD, as the components of TOD primarily revolve around natural language understanding tasks such as NLU, DST, and DPL, while ODD focuses on natural language generation tasks. Specifically, as shown in the bottom part of Figure 3, Meena [30] from Google, Blenderbot [29] from Meta, and EVA [85] from Tsinghua University are all based on encoder-decoder architecture, which demonstrates strong understanding and generation capability simultaneously. In addition, most decoder-only PDMs follow the architecture of GPT2 [25] such as DialoGPT [13] and CDialoGPT [94] for English and Chinese dialogue respectively. Back to this moment, it is observed that most industrial and academic institutions choose to select decoder-only architecture aiming to improve since the scaling law has been widely observed in causal decoders [152].
As the size of the pre-training corpus and the model itself continue to expand, researchers have observed a concurrent improvement in the performance of language models. Throughout this evolution, certain studies have sought to integrate various types of dialogues using meticulously crafted pre-training objectives. The objective is to enable the resulting model to attain enhanced performance in both Task-oriented Dialogue (TOD) and Open-domain Dialogue (ODD) simultaneously [92, 153].
4 Fusion between LLM and DM
Benefiting from scaling law [33], the performance of pre-trained language models (PLM) can be significantly improved by increasing the model size and training data volume. This scaling enables language models to achieve higher levels of understanding and generation capabilities simultaneously. For example, the BERT [24] and GPT series models [11] exemplify this capability, leading to three kinds of fusions during the scaling: 1): fusion within task-oriented dialogue system; 2) fusion between two types of dialogue systems: task-oriented dialogue system and open-domain dialogue system; and 3) fusion between large language model and dialogue model.
4.1 Fusion within TOD
As PLM demonstrates powerful capabilities at different downstream tasks, a lot of work starts to fuse several sub-tasks in TOD (e.g., NLU, DST, DPL, and NLG) and propose a unified end-to-end framework to solve them simultaneously [22, 154, 64, 155, 156, 157], starting from the earliest work NLM (e.g, LSTM, RNN) to the latter PLM (e.g, GPT2, T5), as shown in Table II. By observing Table II, we find that there are several trends accompanied by the shift of used language models: In general, since the appearance of PLM, there is an increasing interest and attention to the end-to-end task-oriented dialogue system as we observe more works appear. Most of them adopt encoder-decoder or decoder-only architectures by formulating each sub-task as a sequence-to-sequence task while leveraging both the understanding and generation capability of language models [158, 157].
Specifically, at the task level in Table II, it is necessary to incorporate the final response generation task (a.k.a, NLG), but the other three sub-tasks are not essential for end-to-end TOD, given that there are multiple methods that do not involve any of the other three tasks. According to the frequency of each task in the different methods, the order relationship is NLG $>$ DST $>$ DPL $>$ NLU. Besides the NLG, DST which tracks the belief state during the conversation becomes more important, since the task-oriented dialogues include many interactions with a specific goal. Furthermore, apart from four key sub-tasks in traditional TOD, we can find more and more additional tasks are optimized with the mentioned key sub-tasks simultaneously. For example, Kulhanek et al. [159] use BC, SAC, and IC to further improve the diversity and quality of responses. Similarly, He et al. [157] use MLM and SRM to learn from large-scale dialog corpora with limited annotations. On the other hand, at the data level, SOLOIST [64], as the first work to investigate the effects of pre-training on heterogeneous dialog corpora, demonstrates great adaptability to accomplish new tasks efficiently. Afterward, lots of studies utilize additional data at the pre-training stage [64, 159, 157] or fine-tuning stage [155, 156], such as He et al. [156] propose GALAXY to explicitly learn dialog policy from limited labeled dialogs and large-scale unlabeled dialog corpora via semi-supervised learning. The only exception is OPERA [160], which proposes a novel dataset OB-MultiWOZ by incorporating question-answering tasks into task-oriented dialogues.
TABLE II: An overview of common E2E Task-oriented Dialogue System, listed in time order.
Model Year Base Model NLU DST DPL NLG Add Tasks* Add Data Evaluation Dataset
NN [22] 2017 LSTM, RNN, CNN Wizard-of-Oz Data
MOSS [154] 2019 GRU CamRest676, LaptopNetwork
MinTL [161] 2020 T5, BART MultiWOZ 2.0
SimpleTOD [162] 2020 GPT2 MultiWOZ 2.0
SOLOIST [64] 2021 GPT2 Cont CamRest676, MultiWOZ 2.0
UBAR [163] 2021 DistilGPT2 MultiWOZ 2.0, 2.1
MTTOD [164] 2021 T5 SP MultiWOZ 2.0, 2.1
AuGPT [159] 2021 GPT2 BC, SAC, IC MultiWOZ 2.0, 2.1
GALAXY [156] 2022 UniLM RS, CR In-Car, MultiWOZ 2.0, 2.1
PPTOD [155] 2022 T5 MultiWOZ 2.0 2.1
OPERA [160] 2022 T5 QA OB-MultiWOZ
SPACE-3 [157] 2022 UniLM MLM,SRM MultiWOZ, CamRest676, In-Car
* Additional task objective abbreviations: Cont-Contrastive Loss, SP–Span Prediction, BC–Belief Consistency, SAC–System Action Classification)
IC–Intent Classification, RS–Response Selection, CR–Consistency Regularization, QA–Question Answering, MLM-Masked Language Modeling
SRM-Semantic Region Modeling
The progression of task-oriented end-to-end dialogue systems has been characterized by notable developments in several key areas. As dialogue systems evolved, encompassing the data and tasks, there emerged a growing emphasis on their evaluation in settings that better simulate real-world scenarios. There are some works that recognize the inherent challenge of realistic settings, such as UBAR [163] is evaluated as a fully end-to-end task-oriented dialogue system where its dialog context has access to user utterances and all content it generated such as belief states, system acts, and system responses. Nonetheless, many studies directly assess the performance of different methods either employing various backbones or incorporating additional data [64, 155, 159]. Given that the choice of pre-trained language models (PLMs) and the pre-training data significantly impact the final performance, the comparison is not that fair, which underscored the necessity for benchmarking in more realistic and fair settings.
The evolution of task-oriented end-to-end dialogue systems has been marked by significant advancements in data, the unification of different tasks, a transition to more realistic settings, and sequence-to-sequence modeling. As the understanding and generation capability of language models keep improving with the scaling of data size and model size, the future holds promise for more versatile and effective task-oriented dialogue systems that can understand and respond to user needs in a wide range of domains and languages [139, 165]. More details can be found in the corresponding survey paper [166].
4.2 Fusion between TOD and ODD
The ongoing pursuit of task-level unification in the field of dialogue systems remains a prominent focus. There are lots of work which have embarked on an exploration of integrated task scenarios in order to achieve a more faithful emulation of human-level conversations, where a diverse array of information is seamlessly integrated [8]. This endeavour extends beyond conventional single dialog types, aiming to integrate ODD and TOD to elevate the authenticity and engagement of dialogue interactions [134, 167, 168, 169, 170]. This kind of mix-type dialogue relax the strong assumption that the user must have a clear and specific goal in TOD, enabling the user to clarify their goals through dynamic interactions [8]. Table III shows the high-level fusion of TOD and ODD.
To fuse different data and tasks between TODs and ODDs, there are different ways according to various types of ODDs: social chit-chat, and knowledge-grounded chit-chat. First of all, several methods attempt to build a unified data schema first while saving some important sub-tasks in TODs and formulating open-domain dialog as end-to-end response generation task [134, 167, 170, 8]. Specifically, both UniDS [134] and PLATO-MT [8] choose to optimize three sub-tasks: belief state generation, system act generation, and response generation, in an end-to-end manner. The data schema usually contains (user input, belief state, DB result, system act, system response). In addition, some of the works consider the transition from TOD to ODD [170] while others model the opposite direction: ODD to TOD [168], both aiming to intertwine these two dialogue modes together seamlessly in the same conversation. Specifically, Tom et al. [170] further feature inter-mode contextual dependency between TOD and ODD, i.e., the dialogue turns from the two modes depending on each other. However, the ODD in these unified dialogue systems is mainly for chit-chat, such as “you’re welcome” and “sounds good to me”, which does not require any external knowledge sources [169].
To incorporate more complicated open-domain situations, another line of work considers knowledge-grounded chit-chat [88, 169], enabling dialog systems to access external knowledge so that they can provide more engaging and knowledgeable conversations and reduce hallucinations simultaneously [171]. Some of them propose to insert additional turns into TOD, where the system needs to respond based on the knowledge from domain FAQs [172] or unstructured documents [173]. These datasets mentioned primarily address the task of identifying conversational turns that necessitate external knowledge for generating responses and selecting the pertinent knowledge for response construction. In contrast, KETOD aims to enhance the authenticity and engagement of dialogues by integrating knowledge-grounded casual conversations into the existing TOD responses [169].
TABLE III: High-level of fusion between TOD and ODD, wherein there are different types of ODD.
Model Task Type of ODD Knowledge sources Training method
Database Web PLMs
Q-TOD [135] TOD End-to-end
Internet-Augmented [88] ODD KG Chit-Chat End-to-end
UniDS [134], FusedChat [170] TOD + ODD Chit-Chat End-to-end
SalesBot [168]
HyKnow [173], KETOD [169] TOD + ODD KG Chit-Chat End-to-end
GODEL [153] TOD + ODD Question Answering, KG Chit-Chat End-to-end
DuClarifyDial [8] TOD + ODD Chit-Chat, Question Answering, KG Chit-Chat End-to-end
MDS [174] TOD + ODD Question Answering End-to-end
OPERA [160] TOD + ODD Question Answering End-to-end
DialogStudio [133] TOD + ODD Dial-Sum, NLU, Conv-Rec, KG Chit-Chat End-to-end
Furthermore, more and more works fuse other tasks or skills into the dialogue system. 1) Skills. Madotto et al. [120] introduce the Adapter-Bot, a dialogue model that uses a fixed backbone conversational model such as DialGPT [13] and triggers on-demand dialogue skills (e.g., emphatic response, weather information, movie recommendation) via different adapters, while Smith et al. [175] also try to blend multiple skills into one conversation session, in which knowledge, emotional and personalizing skills are shown together in the same multi-turn conversation. 2) Tasks. Besides the combination of question-answering and TOD tasks mentioned in § 4.1, DuClarifyDial [8] additionally incorporate chit-chat, knowledge-grounded chit-chat while GODEL [153] specifically target knowledge-grounded goal-directed dialogs by leveraging a new phase of grounded pre-training on open-domain goal-directed corpus. There is the latest work, DialogStudio, which further encompasses data from natural language understanding, conversation recommendation, and dialogue summarization [133]. 3) Knowledge Sources. Most of the previous works either rely on the dataset knowledge for TOD or the Web for ODD [135, 88], and recent works generalize the possible knowledge sources to include databases, the Web, and pre-trained language models at the same time [160]. More analysis can be found in § 5.2.
4.3 Fusion between LLM and DM
TABLE IV: An overview of Dialogue Models based on Large Language Models listed in time order.
Dialogue Models Release Date Base Model # of Params Window size Trainig Tokens Language Preference Source Type
BlenderBot3[89] Aug, 2022 OPT 3B/30B/175B 180B EN Open
ChatGPT Dec, 2022 GPT3.5 [176] 175B 4k/16k EN Closed
Bing Chat Feb, 2023 GPT4 [177] 8k/32k EN Closed
LLaMA-Chat [35] Feb, 2023 LLaMA 7B/13B/33B/65B 2k 1.4T EN Open
Alpaca [178] Mar, 2023 LLaMA 7B 2k 1.4T EN Open
Vicuna [179] Mar, 2023 LLaMA 7B/13B 2k 1.4T EN Open
ChatGLM [180] Mar, 2023 GLM 6B/12B 2k 1T CN Open
Slack Mar, 2023 Claude 200k 2.0T EN Closed
Bard Mar, 2023 PALM 540B EN Closed
Falcon-Chat [181] Jun, 2023 Falcon 7.5B/40B/180B 2k 1.5T/1T/3.5T EN Open
MPT-Chat [182] Jun, 2023 MPT 7B/30B 8k 1T EN Closed
InternLM-Chat [183] Jun, 2023 InternLM 7B/20B 8k 2.3T CN Open
Baichuan-Chat [184], Jun, 2023 Baichuan 7B/13B/53B 4k 1.2T CN Open (7B/13B)
ChatGLM2 [180] Jun, 2023 GLM 6B/12B 32k 1T CN Open
LLaMA2-Chat [185] Jul, 2023 LLaMA2 7B/13B/34B/70B 4k 2.0T EN Open
Atom-Chat Aug, 2023 LLaMA2 7B/13B 4k 1.0T CN Open
Qwen-Chat [186] Aug, 2023 QWen 7B/14B 8k 2.4T CN Open
Baichuan2-Chat [184] Sep, 2023 Baichuan2 7B/13B 4k 2.6T CN Open
ChatGLM3 Oct, 2023 GLM 6B 32k / CN Open
With the emergence of large language models (LLMs), such as ChatGPT and LLaMA [35, 185], the boundaries between task-oriented and chit-chat dialogues have become increasingly indistinct, since LLMs are beginning to demonstrate a preliminary capability for both task-oriented [187] and chit-chat dialogues [188] even under the zero-shot or few-shot settings. Following the release of ChatGPT, numerous large language models are proposed and corresponding dialogue models are finetuned on different-sized dialogue corpus, aiming to achieve better performance on different dialogue tasks, domains or even languages [34, 185]. Hence, an exponential growth in both LLMs and DMs has been observed. To facilitate a more comprehensive illustration and comparison of the advancements, we here choose to focus on the two most widely used languages, namely Chinese and English as shown in Table IV.
LLMs and DMs in English: There are two major series of LLMs: GPT [11, 25, 177] and LLaMA [35, 185]. On one hand, the evolution of the GPT-series models initiated with GPT [11] and has been consistently refined by OpenAI. This improvement involves scaling both the model size and the pre-training data size, including the incorporation of code data [189] and increasing the model parameters to a larger size, notably reaching 175 billion in GPT-3 [176]. The culmination of these enhancements is evident in the success of ChatGPT, achieved through additional fine-tuning and alignment techniques based on the GPT-3.5 model. Initially trained on a dataset of 570GB, ChatGPT supports input lengths of 4k and 16k, and demonstrates remarkable capabilities in human communication, boasting a substantial knowledge base, proficiency in reasoning through mathematical problems, accurate contextual understanding in multi-turn dialogues, alignment with human values to ensure safe use and following diverse user instructions. Subsequently, the introduction of GPT4 [177] further extends the model’s capabilities, enabling the resolution of multi-modal tasks involving both image and text inputs and outputs. It is worthy noting that the maximum input window size is also doubled, which empowers the model to handle long context input. Benefiting from the cooperation with OpenAI, Microsoft release New Bing  [^6] https://blogs.microsoft.com/wp-content/uploads/prod/sites/5/2023/02/The-new-Bing-Our-approach-to-Responsible-AI.pdf, which is a conversational agent based on GPT4, capabling to provide up-to-date information with the assistance of search engine.
On the other hand, unlike the closed-source nature of the GPT series models from OpenAI, Meta has embraced openness by releasing the LLaMA [35] and LLaMA2 [185], which are pre-trained on 1.4T data, offers four parameter configurations: 7B, 13B, 33B, and 65B. This provides a fresh opportunity for an open-source community, with corresponding dialogue models being consistently finetuned based on LLaMA. LLaMA-Chat, for example, finetunes the LLaMA model using instructional methods, aligning with the GPT-3 autoregressive paradigm. In addition, Alpaca [178] is finetuned by Stanford based on LLaMA-7B, using text-davinci-003 API and self-instruct techniques to create a prompt-response instruction dataset of 52K. Vicuna [179], based on LLaMA-13B, used supervised data from ShareGPT  [^7] https://sharegpt.com/, extending sequence length from 512 to 2048 during training. To further improve the window size and human values alignment with the public, LLaMA2 [185] employs grouped-query attention, extending input length from 2k to 4k and training data from 1.4T tokens to 2.0T tokens. Besides these two series, there are other series English LLMs such as Claude [^8] https://claude.ai/, Bard, Falcon, etc. These models typically adhere to an architecture resembling that of GPT-3, showcasing proficiency in handling multi-turn and complex dialogues, with different sizes of input window. Other commercial companies have also adopted similar training methods and subsequently introduced their own conversational large language models, such as MPT-Chat [182] and Falcon-Chat [181].
LLMs and DMs in Chinese: ChatGLM [34, 36], as the first open-source Chinese LLM, which is a bilingual model (Chinese and English) built on the general language model architecture – GLM 130B [36]. The open-sourced 6B version, ChatGLM2-6B, extends the context length from 2k to 32k tokens and employs multi-query attention for more efficient inference. Furthermore, they progress from ChatGLM and ChatGLM2 to ChatGLM3  [^9] https://github.com/THUDM/ChatGLM3/, which incorporates a broader and more diverse corpus of data during its fine-tuning process. Additionally, ChatGLM3 has demonstrated significant improvements in planning and reasoning, akin to language agents [190, 191], thus reflecting substantial improvements in tool learning, code execution, database operation, knowledge graph search and reasoning. There are publicly available LLMs from other companies, such as Baichuan-Chat [184] and Baichuan2-Chat. The latter, Baichuan2-Chat, employs a method similar to Claude and is trained on over 2.6T tokens. These models cater to a wide spectrum of applications, each offering unique advantages stemming from their fine-tuning strategies and underlying architecture. Alibaba releases Qwen-7B-Chat aligns with human intent using a curated dataset, pre-trained on over 2.2T tokens with a context length of 2048. Sensetime publish InternLM-Chat, which is derived from supervised fine-tuning and reinforcement learning with human feedback on InternLM [183], a multilingual model pre-trained on 2.3T tokens.
In summary, as depicted in Table IV, there is a discernible temporal trend towards an increased prevalence of open-sourced LLMs, particularly within the domain of dialogue models. This trend is accompanied by a continuous expansion in the scale of both training data and parameter configurations. The scaling law still works for current LLMs but most of released models is 7B or 13B, due to the computing cost of inferences. Furthermore, there is a noteworthy extension in the window length, denoting an enhanced ability to process and generate longer sequences of text. In the pursuit of improving the utility and safety of dialogue models, there is an evident integration of a higher volume of instruction fine-tuning using conversational data. Additionally, there is a consistent evolution in the multilingual capabilities exhibited by these models.
5 LLM-based Dialogue System
At present, passing the Turing Test and even simulating believable human behavior [192], such as waking up, cooking breakfast, and heading to work, is not a formidable challenge for LLM-based dialogue systems [193, 152]. However, this progress introduces new problems and challenges. First of all, due to unaffordable computing cost during training of LLMs, prompting has emerged as a predominant method for eliciting the reasoning capabilities of LLMs across various tasks [194, 195], including dialogue scenarios [5, 196] (§ 5.1). Secondly, the dialogue system not only interacts with the user but also interacts with external knowledge sources or tools, acquiring necessary intermediate results to resemble the final responses (§ 5.2). These two critical capabilities have become increasingly attractive and crucial for empowering dialogue systems, which can be utilized in complementary to suit the diverse and complicated needs in practice (§ 5.3). Table 4 shows the three major conversational flows for the LLM-based dialogue system.
Figure 4: A novel unified perspective of LLM-based Dialogue System. There are three conversational flows: 1) response based on internal reasoning and dialogue context; 2) response based on external sources and dialogue context; and 3) response based on internal reasoning, external sources, and dialogue context.
5.1 Internal Reasoning
TABLE V: Summary of research works exploring LLMs for various dialogue tasks.
Method(s) Cue(s)† Type Task(s)* LLM(s) Settings Language Domain
Pan et al. [197] S TOD NLU ChatGPT ZS EN General
Michael et al. [198] S DST ChatGPT ZS EN General
IT-LLM [199] S DST, NLG ChatGPT ZS, FS EN General
SGP-TOD [187] S DST, DPL, NLG ChatGPT ZS EN General
GDP-ZERO [200] S DPL ChatGPT FS EN PersuasionForGood
Qin et al. [201] S ODD DS ChatGPT ZS EN General
ProCoT [196] S PD ChatGPT, Vicuna ZS, FS EN General
Fan et al. [202] S DDA ChatGPT ZS, FS CN, EN General
Huynh et al. [203] S DE GPT3, BLOOM FS EN General
Mendonça et al. [204] S DE ChatGPT ZS Multilingual General
Bang et al. [205] S TOD, ODD DRG ChatGPT ZS EN General
CueCoT [5] Per, Psy, E ODD DQA, DRG ChatGPT, ChatGLM ZS, FS CN, EN General
Zhao et al. [206] E DU, DRG ChatGPT ZS EN General
SMILE [207] Psy ODD DRG CN Mental Health
ExTES [208] E ODD DRG LLaMA Finetuning EN General
ChatCounselor [209] Psy DRG Vicuna Finetuning EN Mental Health
CharacterChat [210] Per, M DRG LLaMA Finetuning EN General
Fu et al. [211] E CE, DRG ChatGPT, T5 FS, Finetuning EN General
†S–Semantics, Per–Persona, Psy–Psychology, E–Emotion, M–Memory
* DS–Dialogue Summarization, NLU–Natural Language Understanding, DST–Dialogue State Tracking, PD–Proactive Dialogues, DDA–Dialogue Discourse
Analysis, DE–Dialogue Evaluation, DRG–Dialogue Response Generation, DQA–Dialogue Question Answering, DU–Dialogue Understanding, CE–Causality
Explanation
One crucial aspect is equipping dialogue systems with human-like perception abilities and reasoning abilities. These abilities play a key role in comprehending the internal status of the dialogue context, which stands for any information that can be reasoned and deduced from semantic context during the conversation, benefiting several downstream dialogue tasks: dialogue understanding [206], dialogue state tracking [199, 198, 187], emotional and personalized dialogue generation [5, 208, 211], dialogue summarization [201] and so on. Regarding various prompting methods to reason different aspects of internal status exhibited in the dialogue context, we divide these methods into two categories: 1) semantic perception, which directly extracts useful information from the dialogue context in the required format; 2) semantic reasoning, which reason and deduce some underlying information beneath the dialogue context. Table V shows recent progress in this field.
5.1.1 Semantic Perception
There are some works that attempt to evaluate the performance of LLM at various dialogue understanding tasks such as NLU, DST, and dialogue discourse analysis tasks. For example, Pan et al. [197] and Michael et al. [198] evaluate the capability of ChatGPT to compress the dialogue context into the semantic frame (structured output of NLU) and belief states (structured output of DST) in the context of TOD, respectively. Furthermore, Fan et al. [202] investigate ChatGPT’s performance in three discourse analysis tasks: topic segmentation, discourse relation recognition, and discourse parsing, of varying difficulty levels. The key is to compress the dialogue context, the unstructured text into structured data, to meet the requirement of downstream applications or tasks. On the other hand, another line of work tries to compress the lengthy dialogue context while keeping important semantic information, particularly in dialogue summarization tasks [201].
5.1.2 Semantic Reasoning
Besides semantics information, it is widely acknowledged that dialogue contexts generally convey a lot of information about the user status from a linguistic perspective [212, 213]. Specifically, the linguistic cues underlying dialogue context have been shown to be an effective means of revealing the emotions [214], personality traits [212], psychological characteristics [213], and other relevant information of users [215]. Consequently, recognizing and understanding these cues exhibited in the context of dialogues becomes crucial to comprehend user intentions and status [5]. By doing so, a dialogue system can generate responses that align with the user’s expectations, meet their unique needs and preferences, and ultimately strive towards constructing a human-like conversational experience [5, 7].
Each linguistic cue has garnered considerable scholarly attention, resulting in a wealth of research and a variety of methodological approaches inspired by them. Specifically, Psychology: Liu et al. [209] propose ChatCounselor to provide mental counseling while Qiu et al. [207] utilize ChatGPT to rewrite the QA dataset into multi-turn mental dialogues. Emotion: Fu et al. [211] integrates commonsense-based causality explanation, considering both the user’s perspective (user’s desires and reactions) and the system’s perspective (system’s intentions and reactions). Zhao et al. [206] evaluate the performance of ChatGPT on emotional dialogue in terms of various understanding and generation tasks such as emotion recognition, emotion cause recognition, dialogue act classification (emotion dialogue understanding), empathetic response generation and emotion support generation. In order to navigate limited data and non-standardized training in emotional dialogues, the ExTensible Emotional Support dialogue dataset (ExTES) is curated by utilizing the in-context learning potential of ChatGPT that synthesizes human insights. Persona: Tu et al. [210] present CharacterChat, an extensive conversational system designed to provide social support. CharacterChat is comprised of a conversational model that is guided by personas and memory, alongside an interpersonal matching plugin model. This plugin model is responsible for identifying and assigning the most suitable supporters from the MBTI-1024 Bank to individuals based on their specific personas. All-in-one: Instead of considering different cues in isolation, Wang et al. [5] seamlessly blend them all into one cohesive conversational flow, regarding different aspects of cues as part of user status and prompt the LLMs to reason user status exhibited in the dialogue context, aiming to generate more helpful and acceptable responses for users.
Besides different linguistic cues, it is noteworthy that users may occasionally fail to furnish adequate semantic information within the dialogue context, resulting in ambiguous queries or unreasonable requests. Both of these are considered key aspects of a conversational agent’s proactivity [216, 196]. To trigger the proactivity of LLMs, ProCoT [196] is proposed to augment LLMs with the goal planning capability over descriptive reasoning chains, specifically focusing on three key aspects of proactive dialogues: clarification, target-guided, and non-collaborative dialogues.
5.2 External Sources
To ensure the delivery of up-to-date, informative, trustworthy, and personalized responses, it is necessary to learn and plan how to interact with external different knowledge sources [217]. In this context, the integration of these knowledge sources not only enhances the conversational prowess of the LLM-based dialogue system but also enables them to remain adaptable and aligned with the dynamic nature of the information they handle, ultimately elevating their utility and reliability in real-world scenarios. In this section, we mainly present how LLM-based dialogue systems integrate with these knowledge sources and plan their usage.
  • [leftmargin=*,topsep=0pt,itemsep=0pt,parsep=0pt]
  • Database. TOD mostly relies on one specific database such as the restaurant or flight database to answer the user query, providing the up-to-date and necessary information to accomplish the task [199, 187]. Recently, Hudevceke et al. [199] and Zhang et al. [187] consecutively evaluate the performance of ChatGPT on the E2E TOD. In detail, Hudevceke et al. [199] introduce a novel approach for dialogue state tracking, where belief states are derived from dialog history by utilizing slot descriptions as prompts, subsequently leading to database entry retrieval and response generation. In contrast, SGP-TOD [187] employs slot names and value examples, rather than slot descriptions, as prompts to facilitate frozen LLMs in generating belief states, thereby reducing human effort. In addition, they offer a policy skeleton within the Policy Prompter to guide LLMs in producing responses that comply with the predefined dialog policy, resulting in better performance than IT-LLM [199]. In order to maximize the utilization of databases within dialogue systems, ChatDB [218] employs large-scale models to generate SQL instructions during the dialogue process.
  • Memory. The importance of memory knowledge sources in the context of large language models (LLMs) for open-domain dialogue systems is paramount. These memory sources enable LLMs to retain and access relevant information from previous interactions, facilitating more coherent and contextually aware responses in long-term conversations. Effective memory mechanisms are crucial for a wide range of applications, including personal companion systems, psychological counseling, and secretarial assistance, where sustained interactions and the ability to recall past information are vital for providing a rich and personalized user experience [219, 220, 221]. Particularly, Zhong et al. [220] introduce MemoryBank, a novel memory mechanism inspired by the Ebbinghaus Forgetting Curve theory. MemoryBank allows LLMs to summon and selectively update memories, adapting to a user’s personality over time. It demonstrates the utility of MemoryBank in a long-term AI companion scenario. Wang et al. [221] introduce a recursive memory generation approach, where LLMs are trained to memorize small dialogue contexts and generate new memory as conversations progress, aiming to improve consistency in long-context conversations. Furthermore, Lu et al. [222] present MemoChat, a pipeline for refining instructions to enable LLMs to use self-composed memos to maintain consistency in long-range conversations. It focuses on the iterative “memorization-retrieval-response” process and incorporates tailored instructions. These works collectively contribute to addressing the challenge of long-term memory in LLMs with distinct approaches and innovations.
  • Internet. One key notorious problem of LLMs is the hallucinations [6] [^10] Most works about the hallucination can be found in § 6.2.. Prior studies demonstrate that external document sources can be used to validate the factuality of generated response [223] and thus reduce the hallucinations [171, 6]. In addition, document sources can be used to store up-to-date information, enabling the LLM to sync the latest world knowledge such as New Bing. While responses of LLM-based DS are more likely to include information not present in the relevant document, possibly including the presence of hallucinations, they are preferred by human judges compared with other baselines without using LLM and even human responses [224]. To leverage enhancing the quality and reliability of responses together, WebGLM [225] extends LLMs with web search and retrieval capabilities, demonstrating the efficiency and effectiveness in question-answering tasks, while Semnani et al. [226] mainly leverage documents from Wikipedia to enhance the factuality of LLM-based DS, focusing on providing accurate and verified information with additional fact-checking and refinement steps. Furthermore, there are some works that explore the relationship between multiple documents [227] or queries [228] to solve one complex question requiring multi-hop reasoning.
  • Profile. Salemi et al. [7] first introduce the LaMP benchmark — a novel benchmark for training and evaluating LLMs for producing personalized outputs. LaMP offers a comprehensive evaluation framework with diverse language tasks and multiple entries for each user profile, consisting of three personalized understanding and four personalized generation tasks.
  • Others. Besides these common knowledge sources, there are other knowledge sources in a different organization or structure that also play a key role during response generation such as API [229, 230, 231], and Knowledge Graph [205].
  • All-in-one. Most existing LLM-based DS either focus on a single knowledge source or overlook the relationship between multiple sources of knowledge, which may result in generating inconsistent or even paradoxical responses [232, 233]. To incorporate multiple sources of knowledge and relationships between them, Wang et al. [217] propose SAFARI, a novel framework that takes advantage of the exceptional capabilities of large language models (LLM) in planning, understanding, and incorporating in both supervised and unsupervised settings. Specifically, SAFARI decouples the knowledge grounding into multiple knowledge sources selection and response generation, which allows easy extension to various knowledge sources including the possibility of not using any sources.
5.3 Internal Reasoning with External Sources
By focusing on enhancing the internal status of the dialogue context and allowing seamless interactions with external sources, LLM-based systems can become more sophisticated, effective, and capable of delivering personalized, trustworthy, helpful, up-to-date responses to various user queries and requests [234]. Recently, Tu et al. [210] utilize persona (internal) and profile memory (external) simultaneously to generate personalized responses while Zhang et al. [235] alternatively take advantage of short-term memory (internal) and long-term memory (external). Within the multi-persona collaboration framework TPE (Thinker, Planner, and Executor), Wang et al. [236] leverage the internal status of dialogue context to guide the planning use of different external knowledge sources, regarding different sources as one specific type of conceptual tools. In addition, Sumers et al. [191] propose CoALA, which delineates a language agent that incorporates modular memory components, a structured action space for interactions with both internal memory and external environments, and a generalized decision-making process for action selection.
6 Discussions
6.1 Multi-modality
In the ever-evolving landscape of conversational AI, the integration of multimodal capabilities has emerged as a pivotal advancement that promises to reshape the future of LLM-based dialogue systems [237, 238, 239]. Before the advent of large language models, dialogue systems primarily operated within the confines of text-based interactions, with few works explore the multimodal dialogue systems [240, 241, 242, 243]. However, with the introduction of powerful models such as GPT-4 [177] and their ability to comprehend and generate text and even image at a human-like level, the stage was set for a transformative convergence of text, images, audio, and video in the realm of conversation [237, 238]. In this section, we first briefly review the previous work before the era of large langauge models and then explore the exciting prospects and future directions of multimodal LLM-based dialogue systems, shedding light on their evolution, potential applications, and the dynamic impact they promise to have on a wide range of fields, from customer service and medicine and beyond.
6.1.1 Before the era of LLMs
Before the era of large language models, there were relatively few efforts which try to fuse different modalities seamlessly in the conversational flow. According to different modalities of input and output in dialogue systems, these works can be divided into two categories [^11] https://github.com/ImKeTT/Awesome-Multi-Modal-Dialog: 1) systems with diverse modalities in the input but not in the output [244, 245, 246, 247, 248]; 2) systems with multiple modalities in the output [249, 250, 251, 252]  [^12] This includes systems with the incorporation of multiple modalities in both the input and output.. Predominantly, a significant portion of research endeavors align with the first category, where input comprises text and image modalities, resulting in textual responses pertaining to the image content. A noteworthy milestone in this progression was achieved by Alamri et al. [245] when they pioneered the integration of audio information into the input phase, marking a notable enhancement in the multimodal dialogue system’s capabilities. It is essential to highlight that the majority of efforts falling under the second category primarily involve the generation of output in alternative modalities, such as images, employing a discriminative approach by selecting one from existing sources rather than generating one [240, 243]. Furthermore, some works explore multilingual situations in the context of multimodal dialogue [253, 254].
6.1.2 After the era of LLMs
Since the emergence of LLMs, a notable trend has been the alignment of feature spaces across different modalities within dialogue systems. This alignment effort has taken two distinct directions: 1) the alignment of non-textual modalities with the textual feature space [237, 238, 255, 256, 257]; and 2) approaches that pursue alignment in the opposite direction [258, 259]. These endeavors mark a significant shift in the field, aiming to harmonize the representation and understanding of diverse modalities within the context of LLM-based dialogue systems. In detail, LLaVA represents the inaugural endeavor in visual instruction tuning using data generated by language-only GPT-4, serving as an end-to-end trained large multimodal model that connects a image encoder and LLM for general-purpose visual and language understanding [237]. Despite some preliminary efforts have been made in the development of image-based dialogue systems, some works specifically delve into the relatively uncharted territory of video-based dialogue systems [257, 256]. For example, Muhammad et al. [257] introduce Video-ChatGPT, which is a multimodal model that merges a video-adapted visual encoder with a LLM by simply finetuning a linear layer. Besides that, Zhang et al. [256] further present Video-LLaMA which is capable to understand both visual and auditory content in the video.
In closing, it is crucial to acknowledge that while the evolution of multimodal dialogue systems has been marked by remarkable progress, there still are many challenges and problems. As we forge ahead, we confront an array of practical obstacles that necessitate rigorous research and innovation. The alignment of feature spaces between modalities continues to be a substantial concern, as ensuring a seamless interaction between various data types remains an ongoing challenge. Furthermore, the interpretability and explainability of these complex multimodal systems pose questions regarding their reliability and ethical use. These challenges underscore the need for continued exploration, experimentation, and the development of novel approaches to make multimodal dialogue systems not only powerful but also robust, accessible, and ethically sound, thus solidifying their place in the future of AI-driven conversational technology.
6.2 Hallucination
Language model-based dialogue systems have demonstrated significant potential across a broad range of applications. However, an increasing concern is their observed inclination towards hallucinations [6, 260]. These instances, where system-generated responses defy universally accepted word knowledge, present a substantial challenge to the reliability and effectiveness of these systems in real-world scenarios. As the utilization of these dialogue systems continues to evolve, understanding and mitigating the impact of these hallucinations becomes a critical task for their successful deployment.
6.2.1 Understanding the Causes of Hallucinations
Understanding the underlying causes of hallucinations in language model-based dialogue systems is a crucial first step toward devising effective mitigation strategies. The hallucination phenomenon often results from a complex interplay of several factors. Inadequate training data [261, 262, 263] can play a significant role—when the training data lacks diversity or doesn’t cover certain topics, the model might attempt to fill in the gaps in its knowledge, leading to hallucinated responses. Model overgeneralization [264, 130] is another major contributor, where an language model might overgeneralize from the patterns it has learned during training, leading to responses that may seem plausible but are actually incorrect or nonsensical in a given context. Lastly, the inherent randomness in the generation process can also lead to hallucinations. The stochastic nature of language models [265, 266] means they can sometimes generate incorrect text, particularly when the model is unsure about the most appropriate response. Understanding these underlying causes provides a foundation to explore potential mitigation strategies.
6.2.2 Detecting Hallucinations in Dialogue System
Research on hallucination detection has evolved along four main trajectories. Firstly, human evaluations have been conducted to assess the model output from diverse perspectives [267, 268]. However, these evaluations, while valuable, encounter issues due to their time-consuming nature and the inherent subjectivity of human judgement, leading to scalability and reproducibility challenges. Secondly, datasets have been constructed to evaluate the actuality of open-domain generation with the aid of references [266, 269]. These methods, while offering a more objective evaluation approach, are limited by their dependence on human-labelled references, which can negatively impact their applicability in real-world scenarios and their generalizability to dynamically generated content. Thirdly, self-evaluation methods [195, 270] have been developed to estimate a model’s uncertainty in its output. Despite the simplicity and appeal of these methods, they often lack interpretability and are less effective when dealing with long-form responses. Fourth, some recent studies [261, 271] have applied fact-checking principles to spot factual inaccuracies in the model’s output. These efforts, however, are often constrained by the limitations of external knowledge bases used in the fact-checking process. Finally, an emergent line of research focuses on integrating references or fingerprints into the generated content [272, 273]. This approach, while promising, introduces a layer of traceability and verifiability to the output, enabling users or regulators to trace and verify the origins and accuracy of the content.
6.2.3 Strategies for Mitigating Hallucinations
To mitigate hallucinations in language model-based dialogue systems, researchers have proposed several methods, which can be broadly categorized into five main approaches. One prominent approach is retrieval-augmentation [274], which involves incorporating information from external sources during the generation process to ensure the accuracy of the output. Model editing [275] is another approach adopted by researchers to mitigate hallucinations in language model-based dialogue systems. This involves precisely adjusting the model’s parameters to correct erroneous knowledge inherent in the system. Post-editing [276] is another commonly used approach to mitigate hallucinations in dialogue systems. After the model’s output is generated, post-editing techniques are applied to correct any detected hallucinations. This typically involves rule-based or pattern-based modifications to the generated output, ensuring that the final response is both coherent and accurate. Indeed, some methods aim to mitigate hallucinations by reducing the randomness in language model decoding. For instance, dynamic top-p sampling decoding [266] dynamically adjusts the randomness as the decoding process progresses, thus reducing the element of chance and helping to decrease the incidence of hallucinations. Lastly, better training strategies, such as RLHF [277, 263], have been developed to encourage the model to avoid making factually incorrect statements. By leveraging feedback from humans during the training process, these methods can guide the model towards generating more accurate and reliable responses.
Despite considerable progress in the analysis, detection, and mitigation of hallucinations, they remain a frequent issue in language model-based dialogue systems. Several specific challenges persist. For instance, the wide range of hallucinations, from minor inaccuracies to major factual errors, complicates detection and correction. Furthermore, the lack of large, high-quality datasets for training and evaluation hampers the development of more effective solutions. Balancing creativity and factual consistency is another challenging aspect, as overcorrection could lead to less engaging and overly cautious responses. Lastly, the need for real-time responses in many applications poses a significant challenge, as comprehensive hallucination detection and mitigation can be computationally expensive. Given these challenges, ongoing research and innovation are crucial to improve the effectiveness of these systems in practical applications.
6.3 Safety
Alongside the accelerated evolution of language models and dialogue systems, emerging safety concerns become increasingly crucial. Given the indispensable social impact of current LLM-based dialogue systems, their safety issues have garnered attention not only from the NLP research community [278, 279] but also from industry, philosophers, social scientists [280, 281], policymakers [282], etc. In this subsection, we begin by introducing the safety problem and its distinctive nature in dialogue systems which are directly interacting with human users. Subsequently, we offer a brief overview of existing safeguarding methods and discuss their limitations.
6.3.1 Understanding the Safety Issue
The term safety encompasses a broad spectrum of issues [283], comprising, but not limited to, toxic language [284], social bias [285, 286], problematic moral or values inclination [287, 288], privacy leakage [289], and misleading or risky suggestions [290, 291]. It is noteworthy that the rigorous scope and taxonomy are still an ongoing discussion. Moreover, it is anticipatable that as the system’s capability and application scenarios broaden, the community may encounter novel safety challenges. For example, there is an emergent research trend of investigating prosocial dialogue systems [287], which are sensitive to social factors such as social norms [292, 293], moral stances [294, 288], and human values [295]. This progress stems from the dialogue systems’ expanding capabilities to engage in democratic conversations [280] that involve personal feelings, desires, and social relationships. In light of this, it is crucial to consistently evaluate the performance of dialogue systems to swiftly detect and tackle any new safety concerns that may arise.
6.3.2 Existing Safeguarding Methods
Learning from enormous corpora from various sources, current dialogue systems inevitably learn undesired behavior. Mainstream safeguarding techniques can be categorized into two types: baked-in safety [296] and pipeline safety [278]. Baked-in safety refers to the methods that change the model at the training or decoding stage [297]. The majority of open-source LLMs have incorporated various safety training methods during the training or fine-tuning stages [298, 279], with the RLHF technique [4] being the most effective and widely-discussed option. This method can embed safety into models, however, it is not flexible nor explainable. Pipeline safety does not require modification to models, instead, it uses an add-on safety layer (usually a safety detector) to block unsafe input or output [299]. Nevertheless, this method cannot defend against the malicious use of open-sourced models. Another common and important obstacle in the above methods is the requirement of well-annotated corpora to train classifiers. Recently, there has been a novel research line utilizing the instruction following the ability of LLMs to perform self-critique [300] and self-correct [301]. The LLMs are instructed to serve as the classifier and make post-edition to its responses based on the classification results. Moreover, it is also instructed to self-redteaming [302] to explore high inductive user inputs and enhance the robustness automatically.
6.3.3 Challenges and Discussions
The trade-off between safety and system performance has long been an intricate issue during the development of safe dialogue systems. [296] highlights the challenges of simultaneously maintaining safety and engagement in a dialogue system. Adding to this, Bai et al. [303] provides evidence that there exists a tension between ensuring harmlessness, i.e., safety, and maintaining helpfulness in such systems. This tension, further referred to by researchers as the alignment tax [304] remains an unresolved issue in the field. Another challenge is that the safety standards can change over time, culture, etc. [294] For example, mimic human was an important objective for old-days chatbots [3]. However, nowadays systems are trained to avoid anthropomorphism [305]. Furthermore, the criticism of generalizing the stances and values of a limited group of annotators has not been settled. Considering this, the methods that require a huge amount of training data and extensive training procedure is inflexible. A growing body of literature [288] advocates for steering clear of aligning AI models with specific values, since the human influence during the development lifecycle can inadvertently introduce systematic biases [306, 307]. Instead, scholars argue that shaping AI output through transparent and adaptable guidelines would be a more prudent approach [288]. However, the exploration of this direction remains in its infancy and warrants further investigation to develop a safe dialogue system.
7 Conclusion
This survey presents a comprehensive overview of the dialogue system’s evolution in tandem with advancements in language models, spanning from traditional Statistical Language Models (SLM) to Neural Language Models (NLM), Pre-trained Language Models (PLM), and culminating in the current state of Large Language Models (LLM). The exposition begins by establishing the contextual background of dialogue systems, delineating four distinct stages and a unified definition encompassing various dialogue types. Subsequently, we meticulously detail the progression and transformation of dialogue systems across different language model stages, emphasizing recent strides in LLM-based systems. Concurrently, we succinctly summarize key methodologies characterizing each modeling paradigm at every stage, providing a comprehensive snapshot of current advancements and guiding future directions. Our discourse extends to the application of large language models in addressing emerging challenges such as hallucinations and safety concerns, underscoring the relevance and adaptability of LLMs to evolving problem landscapes. This comprehensive survey not only encapsulates the historical trajectory but also offers insights into the current landscape and potential avenues for future research in dialogue systems.
Acknowledgments
This project is partially supported by HKSAR ITF Project ITT/008/22LP.